OpenStudy (anonymous):

How do you simplify (2xy^2)(3xy)^2 ?

OpenStudy (anonymous):

(2xy^2)(3xy)^2 = (2xy^2)(3xy*3xy) = (2xy^2)(9)(x^2)(y^2) = 18(x)(x^2)(y^2)(y^2) =? // i have the answer, but i just want to see if you know what to do.

OpenStudy (anonymous):

18x^3y^4 ?

OpenStudy (anonymous):

correct